# Refinements of provability and consistency principles for the second incompleteness theorem

### v1

@@ 2.2 Löb’s theorem
Suppose that $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{D2}$ and $\mathbf{D3}^{n}_{n+k}$. Assume $T\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner)\to\varphi$. By the Fixed Point Theorem, we find a sentence $\psi$ satisfying the following equivalence:
$\mathsf{PA}\vdash\psi\leftrightarrow\bigl{(}\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\psi\urcorner)\to\varphi\bigr{)}.$
From $\mathbf{D1}$ and $\mathbf{D2}$, we have
$S\vdash\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)\to\bigl{(}\bigwedge_{2\leq i\leq n+k}\mathrm{Pr}_{T}^{i}(\ulcorner\psi\urcorner)\to\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner)\bigr{)},$
and so
$S\vdash\bigwedge_{1\leq i\leq n+k}\mathrm{Pr}_{T}^{i}(\ulcorner\psi\urcorner)\to\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner).$
Since $\mathrm{Pr}_{T}^{n}(\ulcorner\psi\urcorner)$ is one of the conjuncts of $\bigwedge_{1\leq i\leq n+k}\mathrm{Pr}_{T}^{i}(\ulcorner\psi\urcorner)$, from $\mathbf{D3}^{n}_{n+k}$ and the assumption, we obtain
$T\vdash\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\psi\urcorner)\to\varphi.$
By the fixed point equivalence, $T\vdash\psi$, and hence $S\vdash\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\psi\urcorner)$. We conclude $T\vdash\varphi$. ∎
Corollary 2.14.
For $n,k\geq 1$, $\mathbf{D2}$ $\mathbf{D3}^{n}_{n+k}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{L}}_{T}$.
@@
Relating to this problem, Kogure [13] proved that for each $n,k\geq 1$, there exists a $\Sigma_{1}$ provability predicate whose provability logic is exactly $\mathsf{N}+(\Box^{n}A\to\Box^{n+k}A)$.
@@ 3 Uniform derivability conditions
Here $\ulcorner\varphi(\dot{x})\urcorner$ is a term corresponding to a primitive recursive function calculating the Gödel number of the formula $\varphi(\overline{n})$ from $n$. CB stands for “Converse Barcan”. The implications $\mathbf{\Sigma_{1}C}^{\mathrm{U}}\Rightarrow\mathbf{D3^{\mathrm{U}}}$ and $\mathbf{D1^{\mathrm{U}}}$ $\mathbf{D2^{\mathrm{U}}}\Rightarrow\mathbf{M}^{\mathrm{U}}$ are obvious. The following fact presents more implications.
@@
    ([17, Proposition 2.14.3]) $\mathbf{D2^{\mathrm{U}}}$ $\mathbf{PC^{\mathrm{U}}}\Rightarrow\mathbf{D1^{\mathrm{U}}}$.
@@
Fact 3.2.4 is a refinement of Buchholz’s observation “$\mathbf{D1^{\mathrm{U}}}$ $\mathbf{D2^{\mathrm{U}}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$[4] (see also Rautenberg’s textbook [25]).
@@
    $\mathbf{E}^{\mathrm{U}}$ $\mathbf{PC^{\mathrm{U}}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$.
@@
    $\mathbf{CB_{\exists}}$ $\mathbf{\Delta_{0}C}^{\mathrm{U}}\Rightarrow\mathbf{\Sigma_{1}C}$.
@@
    $\mathbf{CB_{\exists}^{+}}$ $\mathbf{\Delta_{0}C}^{\mathrm{U}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$.
@@
    $\mathbf{E}^{\mathrm{U}}$  $\mathbf{CB^{+}}$.
@@
    $\mathbf{E}^{\mathrm{U}}$  $\mathbf{CB_{\exists}^{+}}$.
@@
    $\mathbf{E}$  $\mathbf{CB}\Rightarrow\mathbf{M}$.
@@
    $\mathbf{E}$  $\mathbf{CB_{\exists}}\Rightarrow\mathbf{M}$.
@@ 3.1 A refinement of Hilbert and Bernays’ G2
$\mathbf{M}$$\mathbf{CB}$ $\mathbf{\Delta_{0}C}^{\mathrm{U}}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$.
@@
$\mathbf{CB}$ $\mathbf{\Delta_{0}C}^{\mathrm{U}}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$.
By combining Theorem 2.8.2 and Proposition 3.4.3, we have “$\mathbf{\Delta_{0}C}^{\mathrm{U}}$ $\mathbf{CB_{\exists}}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{S}}_{T}$”, which may be understood as a dual version of Theorem 3.10.
@@ 3.2 A refinement of the proof of $\mathbf{\Sigma_{1}C}$
Fact 3.2.4 states the implication $\mathbf{M}^{\mathrm{U}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$. By combining this with Proposition 3.5, we have that $\mathbf{E}^{\mathrm{U}}$ $\mathbf{CB_{\exists}^{+}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$. In this subsection, by examining the proof of the latter implication, we show that most parts of the proof require only the assumption $\mathbf{E}^{\mathrm{U}}$, and in particular, we obtain the following new theorem.
@@
$\mathbf{E}^{\mathrm{U}}$ $\mathbf{CB_{\exists}}\Rightarrow\mathbf{\Sigma_{1}C}$.
@@
Then, $\bigvee_{i\leq k}\bigwedge_{j\leq l}(z_{i,j}=t_{i,j}(\vec{x}))$ is exactly $\varphi(\vec{z},\vec{x})\land(\psi(\vec{z},\vec{x})\land z_{k,l}=t_{k,l}(\vec{x}))$. Since $d(t_{k,l}(\vec{x}))=1$, we have that $t_{k,l}(\vec{x})$ is one of 00, $\mathsf{s}(x)$, $x+y$, and $x\times y$.
@@
Case 1: $t_{k,l}(\vec{x})$ is 00.
By substituting 00 for $v$ in (3), since 00 is a numeral, we obtain
@@
Of course, by replacing $\mathbf{CB_{\exists}}$ with $\mathbf{CB_{\exists}^{+}}$ in the above proof of Theorem 3.11, we obtain a proof of “$\mathbf{E}^{\mathrm{U}}$ $\mathbf{CB_{\exists}^{+}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$” which is exactly Fact 3.2.4.
By combining Theorem 3.11, Propositions 3.6 and the fact “$\mathbf{D2}\Leftrightarrow\mathbf{M}$ $\mathbf{C}$”, we obtain the following corollary.
@@
$\mathbf{E}^{\mathrm{U}}$ $\mathbf{CB_{\exists}}$ $\mathbf{C}\Rightarrow\mathbf{D2}$ $\mathbf{D3}$.
@@ 4 Optimality
First, the optimality of “$\mathbf{E}$ $\mathbf{C}$ $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{L}}_{T}$” (Corollary 2.9.1) is guaranteed by the following:
@@
We fix a $\Sigma_{1}$ provability predicate $\exists x\,\mathrm{Proof}_{T}(x,y)$ of $T$, where $\mathrm{Proof}_{T}(x,y)$ is a $\Delta_{1}(\mathsf{PA})$ proof predicate $\mathrm{Proof}_{T}(x,y)$ of $T$ saying that “$x$ is a $T$-proof of a formula $y$”. We may assume $\mathsf{PA}\vdash\mathrm{Proof}_{T}(x,y)\to y\leq x$. We fix a sentence $\xi$ that is not $\Sigma_{1}(T)$, that is, there is no $\Sigma_{1}$ sentence $\sigma$ such that $T\vdash\xi\leftrightarrow\sigma$ (cf. [21, Theorem 1.4]). For each number $m$, let $F_{m}$ be the set of all formulas whose Gödel numbers are less than $m$.
@@
For each natural number $m$, we prove that for all $k$, $(\varphi\in X_{m,k}$ $\xi\notin X_{m})$ implies $T\vdash\varphi$ by induction on $k$. If $(\varphi\in X_{m,0}$ $\xi\notin X_{m})$, then $\varphi$ has a $T$-proof of $\varphi$, and so $T\vdash\varphi$. Suppose that the statement holds for $k$. Assume $(\varphi\in X_{m,k+1}$ $\xi\notin X_{m})$ holds. We distinguish the following four cases:
@@
In the similar way, it is proved that $\mathrm{Pr}_{T}^{\dagger}(x)$ also satisfies $\mathbf{C}$ and $\mathbf{\Sigma_{1}C}$.
@@
We have already confirmed the optimality of “$\mathbf{C}$ $\mathbf{D3}\Rightarrow$ non-$\mathbf{Ros}$” (Theorem 2.10) and “$\mathbf{E}$ $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{S}}_{T}$” (Theorem 2.8).
Next, we consider the optimality of “$\mathbf{\Sigma_{1}C}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{S}}_{T}$” (a consequence of Theorem 2.8).
@@
($(\neg\,\Sigma_{1}(x)\land\mathrm{Pr}_{T}(x))\lor(\Sigma_{1}(x)\land\mathrm{True}_{\Sigma_{1}}(x)),$
@@
At last, we consider “$\mathbf{\Delta_{0}C}^{\mathrm{U}}$ $\mathbf{CB}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$” (Theorem 3.10).
@@
  • $\{\mathbf{\Delta_{0}C}^{\mathrm{U}},\mathbf{CB}\}$ is not sufficient for G2 of $\mathrm{Con}^{\mathrm{S}}_{T}$ by Fact 4.1.3.
@@
    Is there a $\Sigma_{1}$ provability predicate satisfying $\mathbf{\Delta_{0}C}^{\mathrm{U}}$ and $\mathbf{D1^{\mathrm{U}}}$ such that $T\vdash\mathrm{Con}^{\mathrm{H}}_{T}$?

### v2

@@ 2.2 Löb’s theorem
Suppose that $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{D2}$ and $\mathbf{D3}^{n}_{n+k}$. Assume $T\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner)\to\varphi$. Let $\chi$ be any sentence. Since $T\vdash\varphi\land\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\chi\urcorner)\to\varphi$, we have $S\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\land\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\chi\urcorner)\urcorner)\to\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner)$ from $\mathbf{D1}$ and $\mathbf{D2}$. Hence, $T\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\land\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\chi\urcorner)\urcorner)\to\varphi$. Since $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{E}$ and $\mathbf{C}$, it follows from Theorem 2.13 that $T\vdash\varphi$. ∎
@@
Relating to this problem, Kogure [13] proved that for each $n,k\geq 1$, there exists a $\Sigma_{1}$ provability predicate whose provability logic is exactly $\mathsf{N}+(\Box^{n}A\to\Box^{k}A)$.
@@ 3 Uniform derivability conditions
Here $\ulcorner\varphi(\dot{x})\urcorner$ is a term corresponding to a primitive recursive function calculating the Gödel number of the formula $\varphi(\overline{n})$ from $n$. CB stands for “Converse Barcan”. The implications $\mathbf{\Sigma_{1}C}^{\mathrm{U}}\Rightarrow\mathbf{D3^{\mathrm{U}}}$ and $\mathbf{D1^{\mathrm{U}}}$ & $\mathbf{D2^{\mathrm{U}}}\Rightarrow\mathbf{M}^{\mathrm{U}}$ are obvious. The following fact presents more implications.
@@
    ([17, Proposition 2.14.3]) $\mathbf{D2^{\mathrm{U}}}$ & $\mathbf{PC^{\mathrm{U}}}\Rightarrow\mathbf{D1^{\mathrm{U}}}$.
@@
Fact 3.2.4 is a refinement of Buchholz’s observation $\mathbf{D1^{\mathrm{U}}}$ & $\mathbf{D2^{\mathrm{U}}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$[4] (see also Rautenberg’s textbook [25]).
@@
    $\mathbf{E}^{\mathrm{U}}$ & $\mathbf{PC^{\mathrm{U}}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$.
@@
    $\mathbf{CB_{\exists}}$ & $\mathbf{\Delta_{0}C}^{\mathrm{U}}\Rightarrow\mathbf{\Sigma_{1}C}$.
@@
    $\mathbf{CB_{\exists}^{+}}$ & $\mathbf{\Delta_{0}C}^{\mathrm{U}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$.
@@
    $\mathbf{E}^{\mathrm{U}}$ &  $\mathbf{CB^{+}}$.
@@
    $\mathbf{E}^{\mathrm{U}}$ &  $\mathbf{CB_{\exists}^{+}}$.
@@
    $\mathbf{E}$ &  $\mathbf{CB}\Rightarrow\mathbf{M}$.
@@
    $\mathbf{E}$ &  $\mathbf{CB_{\exists}}\Rightarrow\mathbf{M}$.
@@ 3.1 A refinement of Hilbert and Bernays’ G2
$\mathbf{M}$$\mathbf{CB}$ & $\mathbf{\Delta_{0}C}^{\mathrm{U}}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$.
@@
$\mathbf{CB}$ & $\mathbf{\Delta_{0}C}^{\mathrm{U}}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$.
By combining Theorem 2.8.2 and Proposition 3.4.3, we have $\mathbf{CB_{\exists}}$ & $\mathbf{\Delta_{0}C}^{\mathrm{U}}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{S}}_{T}$”, which may be understood as a dual version of Theorem 3.10.
@@ 3.2 A refinement of the proof of $\mathbf{\Sigma_{1}C}$
Fact 3.2.4 states the implication $\mathbf{M}^{\mathrm{U}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$. By combining this with Proposition 3.5, we have that $\mathbf{E}^{\mathrm{U}}$ & $\mathbf{CB_{\exists}^{+}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$. In this subsection, by examining the proof of the latter implication, we show that most parts of the proof require only the assumption $\mathbf{E}^{\mathrm{U}}$, and in particular, we obtain the following new theorem.
@@
$\mathbf{E}^{\mathrm{U}}$ & $\mathbf{CB_{\exists}}\Rightarrow\mathbf{\Sigma_{1}C}$.
@@
Then, $\bigvee_{i\leq k}\bigwedge_{j\leq l}(z_{i,j}=t_{i,j}(\vec{x}))$ is exactly $\varphi(\vec{z},\vec{x})\land(\psi(\vec{z},\vec{x})\land z_{k,l}=t_{k,l}(\vec{x}))$. Since $d(t_{k,l}(\vec{x}))=1$, we have that $t_{k,l}(\vec{x})$ is one of 0, $\mathsf{s}(x)$, $x+y$, and $x\times y$.
@@
Case 1: $t_{k,l}(\vec{x})$ is 0.
By substituting 0 for $v$ in (4), since 0 is a numeral, we obtain
@@
Of course, by replacing $\mathbf{CB_{\exists}}$ with $\mathbf{CB_{\exists}^{+}}$ in the above proof of Theorem 3.11, we obtain a proof of $\mathbf{E}^{\mathrm{U}}$ & $\mathbf{CB_{\exists}^{+}}\Rightarrow\mathbf{\Sigma_{1}C}^{\mathrm{U}}$” which is exactly Fact 3.2.4.
By combining Theorem 3.11, Propositions 3.6 and the fact $\mathbf{D2}\Leftrightarrow\mathbf{M}$ & $\mathbf{C}$”, we obtain the following corollary.
@@
$\mathbf{E}^{\mathrm{U}}$ & $\mathbf{CB_{\exists}}$ & $\mathbf{C}\Rightarrow\mathbf{D2}$ & $\mathbf{D3}$.
@@ 4 Optimality
First, the optimality of $\mathbf{E}$ & $\mathbf{C}$ & $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{L}}_{T}$” (Corollary 2.9.1 and Corollary 2.12) is guaranteed by the following:
@@
We fix a $\Sigma_{1}$ provability predicate $\exists x\,\mathrm{Proof}_{T}(x,y)$ of $T$, where $\mathrm{Proof}_{T}(x,y)$ is a $\Delta_{1}(\mathsf{PA})$ proof predicate $\mathrm{Proof}_{T}(x,y)$ of $T$ saying that “$x$ is a $T$-proof of a formula yy”. We may assume $\mathsf{PA}\vdash\mathrm{Proof}_{T}(x,y)\to y\leq x$. We fix a sentence $\xi$ that is not $\Sigma_{1}(T)$, that is, there is no $\Sigma_{1}$ sentence $\sigma$ such that $T\vdash\xi\leftrightarrow\sigma$ (cf. [21, Theorem 1.4]). For each number $m$, let $F_{m}$ be the set of all formulas whose Gödel numbers are less than $m$.
@@
For each natural number $m$, we prove that for all $k$, $(\varphi\in X_{m,k}$ & $\xi\notin X_{m})$ implies $T\vdash\varphi$ by induction on $k$. If $(\varphi\in X_{m,0}$ & $\xi\notin X_{m})$, then $\varphi$ has a $T$-proof of $\varphi$, and so $T\vdash\varphi$. Suppose that the statement holds for $k$. Assume $(\varphi\in X_{m,k+1}$ & $\xi\notin X_{m})$ holds. We distinguish the following four cases:
@@
In the similar way, it is proved that $\mathrm{Pr}_{T}^{\dagger}(x)$ also satisfies $\mathbf{C}$ and $\mathbf{D3}$.
@@
We have already confirmed the optimality of $\mathbf{C}$ & $\mathbf{D3}\Rightarrow$ non-$\mathbf{Ros}$” (Theorem 2.11) and $\mathbf{E}$ & $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{S}}_{T}$” (Theorem 2.8).
Next, we consider the optimality of $\mathbf{\Sigma_{1}C}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{S}}_{T}$” (a consequence of Theorem 2.8).
@@
$(\neg\,\Sigma_{1}(x)\land\mathrm{Pr}_{T}(x))\lor(\Sigma_{1}(x)\land\mathrm{True}_{\Sigma_{1}}(x)),$
@@
At last, we consider $\mathbf{CB}$ & $\mathbf{\Delta_{0}C}^{\mathrm{U}}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$” (Theorem 3.10).
@@
  • $\{\mathbf{CB},\mathbf{\Delta_{0}C}^{\mathrm{U}}\}$ is not sufficient for G2 of $\mathrm{Con}^{\mathrm{S}}_{T}$ by Fact 4.1.3.
@@
    Is there a $\Sigma_{1}$ provability predicate satisfying $\mathbf{D1^{\mathrm{U}}}$ and $\mathbf{\Delta_{0}C}^{\mathrm{U}}$ such that $T\vdash\mathrm{Con}^{\mathrm{H}}_{T}$?